OpenStudy (anonymous):

A team won 3 of their first 10 games. If they win the next n games, they will have won 50% of their games. What is the value of n?

OpenStudy (anonymous):

(n + 3)/(n + 10) = 1/2 n + 3 = 1/2(n + 10) n + 3 = n/2 + 5 n/2 = 2 n = 4